# Overlap Gap and Computational Thresholds in the Square Wave Perceptron

### v2

@@ 3.2.2 Failure of the Annealed and RS Ansatz for $\delta\neq 0$
As discussed in Section 2, the presence of $m$-OGP implies $m^{\prime}$-OGP for $m<m^{\prime}$. Therefore, for $m<m^{\prime}$, $\alpha_{\mathrm{\mathrm{OGP}}}(m)\leq\alpha_{\mathrm{OGP}}(m^{\prime})$. However this ordering is not to be respected at the level of the annealed estimates of the thresholds for $\delta\neq 0$, as $\alpha_{\mathrm{OGP}}^{\mathrm{ann}}(m,\delta)$ is a non-monotone function of the number of clones $m$, see Fig. 4(a). This was also observed for $\text{SBP}_{\kappa}$ in [6] for finite $\kappa$. Moreover, in Fig. 3(a) one can also observe that even taking small values of $m$, the corresponding $\alpha_{\mathrm{ogp}}$ intercepts the SAT/UNSAT transition at a certain value of $\delta$. This is clearly unphysical since in the ABP, that corresponds to $\delta\to\infty$, it has already been observed that an hard algorithmic phase exists. All those inconsistencies are due to the annealed computation of the $m-$OGP thresholds, which only give an upper bound of the true value. We have therefore resorted to a replica symmetric calculation of $\alpha_{\mathrm{OGP}}$, see F.
@@
We show $\alpha^{\mathrm{rs}}_{\mathrm{OGP}}(\delta)$ for $m=2,3,4$ in Fig. 3(b). For $\delta\to 0$, the RS estimation coincides with the annealed one, as expected. However, for intermediate and large values of $\delta$ the RS estimation significantly deviates from the annealed one. In particular, for $\delta\to\infty$ and $m=4$ one obtains $\alpha^{\mathrm{rs}}_{\mathrm{OGP}}\simeq 0.7843$ which is close to what was estimated in [16, 28]. However, in Fig. 4(b) we show that for fixed $\delta\neq 0$, even the RS estimate of the OGP threshold suffers the same problem of the annealed approximation, i.e. $\alpha_{\mathrm{OGP}}^{\mathrm{rs}}(m)$ is still a non-monotonic function of $m$. This is also signaled by the fact that as soon as $\alpha_{\mathrm{OGP}}^{\mathrm{rs}}(m)$ starts increasing with $m$, the external entropy becomes negative (see Fig. 12).
@@
Figure 3: (a): Storage capacity $\alpha_{sat}$, and $m$-OGP thresholds of the SWP as a function of $\delta$. In particular, for each $\delta$, starting from top to bottom, the colored curves represent the $m$-OGP thresholds for $m=2,\dots,m^{\star}(\delta)$, where $m^{\star}(\delta)=\min{\{15,\text{argmin}_{m}\{\alpha_{\mathrm{OGP}}(m)\}\}}$. The dots at $\delta=0$ correspond to $1/m$. The black curve at the bottom represents $\min_{m}\widetilde{\alpha}^{ann}_{OGP}(m,\delta)$ (see Sec. 3.2). Inset: for large $\delta$, $\alpha_{sat}$ reaches a plateau (dashed line) corresponding to the storage capacity of the asymmetric binary perceptron $\alpha_{c}^{ABP}\approx 0.833$ (see [29]). (b): Same as left panel, with the addition of he dashed lines, which represent the replica symmetric corrections to the annealed $m$-OGP estimates for $m=2,3,4$. For small values of $\delta$ the RS predictions collapse on the annealed curves. For large $\delta$, the RS threshold deviate from the annealed ones.
@@
Figure 4: (a): First-moment upper bound of $\alpha_{\mathrm{OGP}}$ in the SWP for different values of $\delta$ in the storage setting. From top to bottom $\delta=1.5,1.2,1.1,1$. The true $\alpha_{\mathrm{OGP}}$ must be a non-increasing function of $m$. The fact that the first moment estimates are non-monotone, implies that the annealed computation cannot be exact. Note that the minimum shifts to larger values of $m$ upon decreasing $\delta$. (b): Annealed (upper curve) and replica symmetric (lower curve) estimates of the $\alpha_{OGP}(m)$ thresholds as a function of $m$ for $\delta=1.5$. The replica symmetric ansatz does not cure the unphysical non-monotonicity of the annealed estimates of the thresholds.
@@ 3.3 Algorithmic Attack
In a series of works [48, 49, 50, 51, 16, 41], it has been shown that variations of message passing algorithms derived from the cavity method provide powerful solvers for ABP and other CSPs. In this section, we study the behavior of one of these algorithms, called reinforced approximate message passing, (RAMP). In RAMP each weight $w_{i}$ of the perceptron is associated with a “magnetization” $m_{i}$, with $-1\leq m_{i}\leq 1$. Magnetizations are updated according to local rules that depend on an external parameter $r$, called reinforcement rate, which has to be adjusted heuristically. We call iteration the set of $O(PN)$ operations needed to update the magnetizations of all sites. After each iteration, the binary vector $\widetilde{\boldsymbol{w}}$, with elements $\widetilde{w}_{i}=\text{sgn}(m_{i})$ is computed and checked to see if it is a solution to the problem. We refer the reader to G for the details of the algorithm. In Fig. 5(a) we show that the number of iterations $T_{sol}$ required to find a solution exhibits a critical behavior, namely it diverges in correspondence to an algorithmic threshold $\alpha_{alg}(\delta)$ that depends on $\delta$. In Fig. 5(b) we show the difference $\Delta\alpha=\min_{m}\alpha_{OGP}^{rs}(m,\delta)-\alpha_{alg}(\delta)$, between the best replica-symmetric estimate $\min_{m}\alpha_{OGP}^{rs}(m,\delta)$ of the $m$-OGP threshold and an extrapolation of the algorithmic threshold $\alpha_{alg}(\delta)$, as a function of $\delta$. The shift $\Delta\alpha$ displays a non-monotone behavior. For the largest value of $\delta$ in the dataset, $\delta=1.5$, the shift $\Delta\alpha$ is compatible with zero. This is also the case in the limit of $\delta\rightarrow\infty$, where [16] showed agreement between the estimates of the algorithmic threshold of solvers based on the cavity method and a RS local entropy computation. An intermediate region for $\delta<1.5$ displays positive shifts. However, as we have discussed in Section 3.2.2, the RS computation is not exact. Breaking the replica symmetry might be needed to understand this gap.
@@
Figure 5: (a): average number of iterations $T_{sol}$ as a function of $\alpha$, for different values of $\delta$. From left to right $\delta=0.6,1,1.1,1.2,1.5$. The behavior of $T_{sol}$ is compatible with $T_{sol}\sim\mathrm{e}^{\frac{a}{\alpha_{alg}(\delta)-\alpha}}$, implying that for small $\alpha_{alg}(\delta)-\alpha$, $\log^{-1}{(r_{opt})}\propto\alpha-\alpha_{alg}(\delta)$. This allows to estimate the algorithmic thresholds $\alpha_{alg}(\delta)$ (see the inset of Fig. 5(b)). Data are obtained averaging over $80$ samples of size $N=4000$. (b): difference $\Delta\alpha(\delta)$ between the optimum value of the RS estimate of the $m$-OGP threshold, namely $\min_{m}\alpha_{OGP}^{rs}(m)$, and the algorithmic thresholds estimated in figure (a). Inset: $\log^{-1}{(r_{opt})}$ as a function of $\alpha$ for different $\delta$. From left to right $\delta=0.6,1,1.1,1.2,1.5$. Dashed lines are linear fit of the last points.
@@ 4 The Teacher-Student Problem
In the planted setting, it is natural to distinguish between two algorithmic problems: optimization, which asks to find a generic solution, and learning, or recovery, which asks to find the teacher. Note that in this case, for all values of $\alpha$, there is by definition at least one solution, the teacher $\boldsymbol{w}^{T}$. Depending on the activation function, one finds a critical value $\alpha_{T}(\varphi)$, the “teacher threshold”, s.t. for $\alpha\geq\alpha_{T}(\varphi)$, $\boldsymbol{w}^{T}$ is the unique solution with high probability, while for $\alpha<\alpha_{T}(\varphi)$ an exponential number of solution exist. Below $\alpha_{T}(\varphi)$ the learning task is impossible, since the teacher is indistinguishable from the other solutions. Above $\alpha_{T}(\varphi)$ optimization and learning are the same problem.
In Section 4.1 we compute $\alpha_{T}(\varphi)$ for the SWP. In Section 4.2 we study the presence of an overlap-gap for $\alpha<\alpha_{T}(\varphi)$. In Section 4.3 we discuss the recovery problem.
@@ 4.1 The Teacher Threshold
$\mathcal{N}_{1}=\sum_{\boldsymbol{w}}\prod_{\mu=1}^{P}\Theta\left[\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}^{T}\cdot\boldsymbol{x}^{\mu}\right)\,\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}\cdot\boldsymbol{x}^{\mu}\right)\right]\,.$ (26)
Using an interpolation method, [52] proves that if the patterns $\boldsymbol{x}^{\mu}$ have i.i.d. elements, and the activation is continuous almost everywhere, the replica-symmetric free entropy for the teacher-student problem is exact. Similarly to the storage case, the replica-symmetric quenched free entropy of the teacher-student problem can be obtained as the limit $m\rightarrow 0$ of the annealed clonated free-entropy, see B.1:
@@
Note that in this case, the optimization is over two parameters, the overlaps $p$ and $q$. However one finds that if the problem is self-averaging, the parameter to fix is just one, since on the fixed point $p=q$. This is a consequence of the so-called Nishimori identities [53, 54, 55, 56]. One can show that for any $\alpha$, the teacher (i.e. $p=1$) is always a maximum with zero entropy. For $\alpha<\alpha_{T}$, it is only a local maximum, as there exists another fixed point $p<1$ with positive entropy. For $\alpha>\alpha_{T}$ the point at $p=1$ becomes a global maximum of the free entropy. In Fig. 2, we derive $\alpha_{T}$ as a function of $\delta$ for the SWP. Analogously to the storage case, in the small-$\delta$ region, $\alpha_{T}$ saturates the lower bound $\alpha_{T}\geq 1$. An upper-bound is provided by the annealed computation of the free entropy (see B.2).
@@ 4.2.1 The Small $\delta$ Limit
for some coefficients $c_{1},\dots,c_{m}$. Note that there cannot be a zero coefficient in front of $u_{0}$ since the rows corresponding to the students are linearly independent. Note also that since the $u$’s have unitary variance:
@@
The request that $\Sigma$ is a covariance matrix implies a further constraint. Indeed, it should be always possible to write the elements of $\Sigma$ as $\Sigma_{ab}=\frac{1}{N}\sum_{i}\tau_{a}^{i}\,\tau_{b}^{i}$, for some $\tau_{a}^{i}=\pm 1$. Therefore if $\tilde{\boldsymbol{u}}$ is an eigenvector of $\Sigma$ associated with a zero eigenvalue, one has:
@@
Suppose now that there are at least two non-zero coefficients. Squaring Eq. 35, and subtracting Eq. 33 we get the following condition:
@@
However this equation is not solvable by any $\tau_{1},\dots,\tau_{m}\in\{\pm 1\}^{m}$ for any $m$ unless $q=1$. Indeed the right-hand side of Eq. (41) is always greater than or equal to $m$ (with equality attained for $q=1$), whereas the left-hand side can only be an integer in the range $-m,-m+2,\dots,m-2,m$. This means that the only correlation matrices satisfying Eq. (39) are those produced by configurations that differ from the all ones vector by at most $o(m)$ bits. Therefore if $q<1$ it is not possible to choose more than one non-zero $c_{a}$, and in the limit $\delta\to 0$ the only possibility is that the energetic term reduces to (30). By the same argument exposed in the storage case in Section 3.2.1, this implies that for $\delta\to 0$ the $m$-OGP threshold is given by $1/m$ as in Eq. (25) and vanishes in the large $m$ limit.
Figure 6: (a): Starting from the top right, the dashed line represents the hard-easy transition (see Sec. 4.3), above which AMP is able to fully reconstruct the teacher, as a function of $\delta$. Note that for $\delta\rightarrow\infty$ the hard-easy tends to $\alpha_{r}\approx 1.49$, which is the spinodal transition of the planted ABP. The black continuous line represents the teacher threshold $\alpha_{T}$, above which the teacher is the unique solution, as a function of $\delta$. The colored lines are the annealed upper bounds of the $m$-OGP, as a function of $\delta$. In particular, going from top to bottom, $m=2,\dots,m^{\star}(\delta)$, where $m^{\star}(\delta)=\min{\{15,\text{argmin}_{m}\{\alpha_{\mathrm{OGP}}(m)\}\}}$. The dots at $\delta=0$ correspond to $1/m$. (b): The dashed line represents the replica symmetric corrections to the annealed $m$-OGP estimates for $m=2,3,4$. From small values of $\delta$ the RS predictions collapse on the annealed curves. For large $\delta$, the RS threshold deviate from the annealed ones.

### v4

@@ 3.2.2 Failure of the Annealed and RS Ansatz for $\delta\neq 0$
As discussed in Section˜2, the presence of $m$-OGP implies $m^{\prime}$-OGP for $m<m^{\prime}$. Therefore, for $m<m^{\prime}$, $\alpha_{\mathrm{\mathrm{OGP}}}(m)\leq\alpha_{\mathrm{OGP}}(m^{\prime})$. However this ordering is not to be respected at the level of the annealed estimates of the thresholds for $\delta\neq 0$, as $\alpha_{\mathrm{OGP}}^{\mathrm{ann}}(m,\delta)$ is a non-monotone function of the number of clones $m$, see Fig.˜4(a). This was also observed for $\text{SBP}_{\kappa}$ in [6] for finite $\kappa$. Moreover, in Fig.˜3(a) one can also observe that even taking small values of $m$, the corresponding $\alpha_{\mathrm{ogp}}$ intercepts the SAT/UNSAT transition at a certain value of $\delta$. This is clearly unphysical since in the ABP, that corresponds to $\delta\to\infty$, it has already been observed that an hard algorithmic phase exists. All those inconsistencies are due to the annealed computation of the $m$-OGP thresholds, which only give an upper bound of the true value. We have therefore resorted to a replica symmetric calculation of $\alpha_{\mathrm{OGP}}$, see H.
@@
We show $\alpha^{\mathrm{rs}}_{\mathrm{OGP}}(\delta)$ for $m=2,3,4$ in Fig.˜3(b). For $\delta\to 0$, the RS estimation coincides with the annealed one, as expected. However, for intermediate and large values of $\delta$ the RS estimation significantly deviates from the annealed one. In particular, for $\delta\to\infty$ and $m=4$ one obtains $\alpha^{\mathrm{rs}}_{\mathrm{OGP}}\simeq 0.7843$ which is close to what was estimated in [16, 35]. However, in Fig.˜4(b) we show that for fixed $\delta\neq 0$, even the RS estimate of the OGP threshold suffers the same problem of the annealed approximation, i.e. $\alpha_{\mathrm{OGP}}^{\mathrm{rs}}(m)$ is still a non-monotonic function of $m$. This is also signaled by the fact that as soon as $\alpha_{\mathrm{OGP}}^{\mathrm{rs}}(m)$ starts increasing with $m$, the external entropy becomes negative (see Fig.˜13).
@@
Figure 3: (a): Storage capacity $\alpha_{sat}$, and annealed $m$-OGP thresholds of the SWP as a function of $\delta$. In particular, for each $\delta$, starting from top to bottom, the colored curves represent the annealed $m$-OGP thresholds for $m=2,\dots,m^{\star}(\delta)$, where $m^{\star}(\delta)=\min{\{15,\text{argmin}_{m}\{\alpha_{\mathrm{OGP}}(m)\}\}}$. The dots at $\delta=0$ correspond to $1/m$. The black curve at the bottom represents $\min_{m}\widetilde{\alpha}^{ann}_{OGP}(m,\delta)$ (see Sec. 3.2). Inset: for large $\delta$, $\alpha_{c}$ reaches a plateau (dashed line) corresponding to the storage capacity of the asymmetric binary perceptron $\alpha_{c}^{ABP}\approx 0.833$ (see [36]). (b): Same as left panel, with the addition of the dashed lines, which represent the replica symmetric corrections to the annealed $m$-OGP estimates for $m=2,3,4$. For small values of $\delta$ the RS predictions collapse on the annealed curves. For large $\delta$, the RS threshold deviate from the annealed ones.
@@
Figure 4: (a): First-moment upper bound of $\alpha_{\mathrm{OGP}}$ in the SWP for different values of $\delta$ in the storage setting. From top to bottom $\delta=1.5,1.2,1.1,1$. The true $\alpha_{\mathrm{OGP}}$ must be a non-increasing function of $m$. The fact that the first moment estimates are non-monotone, implies that the annealed computation cannot be exact. Note that the minimum shifts to larger values of $m$ upon decreasing $\delta$. (b): Annealed (upper curve) and replica symmetric (lower dashed curve) estimates of the $\alpha_{OGP}(m)$ thresholds as a function of $m$ for $\delta=1.5$. The replica symmetric ansatz does not cure the unphysical non-monotonicity of the annealed estimates of the thresholds.
@@ 3.3 Algorithmic Attack
In a series of works [34, 33, 54, 55, 16, 47], it has been shown that variations of message passing algorithms derived from the cavity method provide powerful solvers for ABP and other CSPs. In this section, we study the behavior of one of these algorithms, called reinforced approximate message passing, (RAMP). In RAMP each weight $w_{i}$ of the perceptron is associated with a “magnetization” $m_{i}$, with $1\leq m_{i}\leq 1$. Magnetizations are updated according to local rules that depend on an external parameter $r$, called reinforcement rate, which has to be adjusted heuristically. We call iteration the set of $O(PN)$ operations needed to update the magnetizations of all sites. After each iteration, the binary vector $\widetilde{\boldsymbol{w}}$, with elements $\widetilde{w}_{i}=\text{sgn}(m_{i})$ is computed and checked to see if it is a solution to the problem. We refer the reader to I for the details of the algorithm. In Fig.˜5(a) we show that the number of iterations $T_{sol}$ required to find a solution exhibits a critical behavior, namely it diverges in correspondence to an algorithmic threshold $\alpha_{alg}(\delta)$ that depends on $\delta$. In Fig.˜5(b) we show the difference $\Delta\alpha=\min_{m}\alpha_{OGP}^{rs}(m,\delta)-\alpha_{alg}(\delta)$, between the best replica-symmetric estimate $\min_{m}\alpha_{OGP}^{rs}(m,\delta)$ of the $m$-OGP threshold and an extrapolation of the algorithmic threshold $\alpha_{alg}(\delta)$, as a function of $\delta$. The shift $\Delta\alpha$ displays a non-monotone behavior. For the largest value of $\delta$ in the dataset, $\delta=1.5$, the shift $\Delta\alpha$ is compatible with zero. This is also the case in the limit of $\delta\rightarrow\infty$, where [16] showed agreement between the estimates of the algorithmic threshold of solvers based on the cavity method and a RS local entropy computation. An intermediate region for $\delta<1.5$ displays positive shifts. However, as we have discussed in Section˜3.2.2, the RS computation is not exact. Breaking the replica symmetry might be needed to understand this gap.
@@
Figure 5: (a): average number of iterations $T_{sol}$ as a function of $\alpha$, for different values of $\delta$. From right to left $\delta=1.5,1.2,1.1,1,0.6$. The behavior of $T_{sol}$ is compatible with $T_{sol}\sim\mathrm{e}^{\frac{a}{\alpha_{alg}(\delta)-\alpha}}$, implying that for small $\alpha_{alg}(\delta)-\alpha$, $\log^{-1}{(r_{opt})}\propto\alpha-\alpha_{alg}(\delta)$. This allows to estimate the algorithmic thresholds $\alpha_{alg}(\delta)$ (see the inset of Fig.˜5(b)). Data are obtained averaging over 8080 samples of size $N=4000$. (b): difference $\Delta\alpha(\delta)$ between the optimum value of the RS estimate of the $m$-OGP threshold, namely $\min_{m}\alpha_{OGP}^{rs}(m)$, and the algorithmic thresholds estimated in figure (a). Inset: $\log^{-1}{(T_{sol})}$ as a function of $\alpha$ for different $\delta$. From right to left $\delta=1.5,1.2,1.1,1,0.6$. Dashed lines are linear fit of the last points.
@@ 4 The Teacher-Student Problem
In the planted setting, it is natural to distinguish between two algorithmic problems: optimization, which asks to find a generic solution, and learning, or recovery, which asks to find the teacher. Note that in this case, for all values of $\alpha$, there is by definition at least one solution, the teacher $\boldsymbol{w}^{*}$. Depending on the activation function, one finds a critical value $\alpha_{T}(\varphi)$, the “teacher threshold”, s.t. for $\alpha\geq\alpha_{T}(\varphi)$, $\boldsymbol{w}^{*}$ is the unique solution with high probability, while for $\alpha<\alpha_{T}(\varphi)$ an exponential number of solution exist. Below $\alpha_{T}(\varphi)$ the learning task is impossible, since the teacher is indistinguishable from the other solutions. Above $\alpha_{T}(\varphi)$ optimization and learning are the same problem.
In Section˜4.1 we compute $\alpha_{T}(\varphi)$ for the SWP. In Section˜4.2 we study the presence of an overlap-gap for $\alpha<\alpha_{T}(\varphi)$. In Section˜4.3, we discuss the recovery problem.
@@ 4.1 The Teacher Threshold
$\mathcal{N}_{1}=\sum_{\boldsymbol{w}}\prod_{\mu=1}^{P}\Theta\left[\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}^{*}\cdot\boldsymbol{x}^{\mu}\right)\,\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}\cdot\boldsymbol{x}^{\mu}\right)\right]\,.$ (26)
Using an interpolation method, [56] proves that if the patterns $\boldsymbol{x}^{\mu}$ have i.i.d. elements, and the activation is continuous almost everywhere, the replica-symmetric free entropy for the teacher-student problem is exact. Similarly to the storage case, the replica-symmetric quenched free entropy of the teacher-student problem can be obtained as the limit $m\rightarrow 0$ of the annealed clonated free-entropy, see C.1:
@@
Note that in this case, the optimization is over two parameters, the overlaps $p$ and $q$. However, one finds that if the problem is self-averaging, the parameter to fix is just one, since on the fixed point $p=q$. This is a consequence of the so-called Nishimori identities [57, 58, 59, 60]. One can show that for any $\alpha$, the teacher (i.e. $p=1$) is always a maximum with zero entropy. For $\alpha<\alpha_{T}$, it is only a local maximum, as there exists another fixed point $p<1$ with positive entropy. For $\alpha>\alpha_{T}$, the point at $p=1$ becomes a global maximum of the free entropy. In Fig.˜2, we derive $\alpha_{T}$ as a function of $\delta$ for the SWP. Analogously to the storage case, in the small-$\delta$ region, $\alpha_{T}$ saturates the lower bound $\alpha_{T}\geq 1$. An upper-bound is provided by the annealed computation of the free entropy (see C.2).
@@ 4.2.1 The Small $\delta$ Limit
This follows from the fact that the theta functions tessellate the (m+1)(m{+}1)-dimensional integration space into boxes of side $\delta$, alternating between values $1$ and 0. Since the Gaussian measure is regular, in the limit $\delta\to 0$ the integral reduces to the average value of the boxes, which equals $1/2$.
@@
for some coefficients $c_{1},\dots,c_{m}$. Note that there cannot be a zero coefficient in front of $u_{0}$ since the rows corresponding to the students are linearly independent. Note also that the $u$’s have unitary variance:
@@
The request that $\Sigma$ is a covariance matrix implies a further constraint. Indeed, it must be possible to write the elements of $\Sigma$ as $\Sigma_{ab}=\frac{1}{N}\sum_{i}\tau_{a}^{i}\,\tau_{b}^{i}$, for some $\tau_{a}^{i}=\pm 1$. This constraint is embodied in the entropic contribution to the free entropy, which can be written as (see Eq.˜70)
$G_{S}^{(m)}(q,p_{1},\dots,p_{m})=\frac{1}{mN}\log{\frac{1}{2^{N}}\sum_{\boldsymbol{w}_{0}\cdots\boldsymbol{w}_{m}}\prod_{a<b}\delta\left(q-\frac{\boldsymbol{w}_{a}\cdot\boldsymbol{w}_{b}}{N}\right)\,\prod_{a}\delta\left(p_{a}-\frac{\boldsymbol{w}_{0}\cdot\boldsymbol{w}_{a}}{N}\right)}$ (34)
Any $\Sigma$ not fulfilling the above constraint would lead to an infinitely negative $G_{S}$. As we will be interested in the $N\to\infty$ limit, the quantized nature of the value of $q,p_{1},$... is lost, but other implications of the constraint remain strong even in this limit [^3] We note here that the requirement that $\Sigma$ can be written as $\Sigma_{ab}=\frac{1}{N}\sum_{i}\tau_{a}^{i}\,\tau_{b}^{i}$ is stronger than positive definitiveness. For example it is easy to check that the matrix with $m=2$ $\begin{pmatrix}1&p_{1}&p_{2}\\ p_{1}&1&q\\ p_{2}&q&1\end{pmatrix}$ is not acceptable whenever $p_{2}\leq q+p_{1}-1$ (as it violates triangular inequality), but can be positive definite.. Namely, if $\tilde{\boldsymbol{u}}$ is an eigenvector of $\Sigma$ associated with a zero eigenvalue, one has:
@@
Hence, there must exist at least one binary sequence $\tau_{1},\dots,\tau_{m}\in\{\pm 1\}^{m}$ s.t.
$\sum_{a}\tau_{a}^{i}\,\tilde{u}_{a}=0.$ (36)
@@
Suppose now that there are at least two non-zero coefficients. Squaring Eq.˜37, and subtracting Eq.˜33 we get the following condition:
@@
However this equation is not solvable by any $\tau_{1},\dots,\tau_{m}\in\{\pm 1\}^{m}$ for any $m$ unless $q=1$. Indeed the right-hand side of Eq. (43) is always greater than or equal to $m$ (with equality attained for $q=1$), whereas the left-hand side can only be an integer in the range $m,-m+2,\dots,m-2,m$. This means that the only correlation matrices satisfying Eq. (41) are those produced by configurations that differ from the all ones vector by at most $o(m)$ bits. Therefore if $q<1$ it is not possible to choose more than one non-zero $c_{a}$, and in the limit $\delta\to 0$ the only possibility is that the energetic term reduces to (30). By the same argument exposed in the storage case in Section˜3.2.1, this implies that for $\delta\to 0$ the $m$-OGP threshold is given by $1/m$ as in Eq. (25) and vanishes in the large $m$ limit.
Figure 6: (a): Starting from the top right, the dashed line represents the hard-easy transition (see Sec. 4.3), above which AMP is able to fully reconstruct the teacher, as a function of $\delta$. For $\delta\rightarrow\infty$ the hard-easy tends to $\alpha_{r}\approx 1.49$, which is the spinodal transition of the planted ABP. Moreover notice how this line is non-monotonic in $\delta$, for an intuitive explanation see E. The black continuous line represents the teacher threshold $\alpha_{T}$, above which the teacher is the unique solution, as a function of $\delta$. The colored lines are the annealed upper bounds of the $m$-OGP, as a function of $\delta$. In particular, going from top to bottom, $m=2,\dots,m^{\star}(\delta)$, where $m^{\star}(\delta)=\min{\{15,\text{argmin}_{m}\{\alpha_{\mathrm{OGP}}(m)\}\}}$. The dots at $\delta=0$ correspond to $1/m$. (b): The dashed line represents the replica symmetric corrections to the annealed $m$-OGP estimates for $m=2,3,4$. From small values of $\delta$ the RS predictions collapse on the annealed curves. For large $\delta$, the RS threshold deviate from the annealed ones.